# Towards Comprehensive Vietnamese Retrieval-Augmented Generation and Large Language Models

### 03 Mar 2024 v1

@@ 2 Main contributions
  • A massive Vietnamese NewsCorpus dataset of around 32M articles, a substantial 53 GB in size, rigorously cleaned, deduplicated, and formatted specifically for the continual pre-training of LLMs.
@@
  • An additional large-scale Vietnamese NewsCategory dataset in a “text-category” format, specifically designed for text classification task.
  • Vietnamse Alpaca datasets, tailored for supervised fine-tuning LLMs.
@@
  • We also offer two base models, Vietnamese LLaMA2-7b, which have been further pretrained on an expansive corpus of Vietnamese text, 40 GB and 120 GB respectively, derived from LLaMA2, marking a significant advancement in the understanding and generation of the Vietnamese language.
@@ 3.2 Vietnamese NewsSapo dataset
The Vietnamese NewsSapo dataset was constructed to training of sentence/passage embeddings. Our dataset is structured in a ”title-abstract-contents” format, where each news article is represented by a tuple of (title, abstract, content). The content is the main text body of the article and has been processed to remove images, videos, and other non-textual elements. The dataset contains 31,728,183 triples.
@@ 3.3 Vietnamese NewsCategory dataset
The Vietnamese NewsCategory dataset is constructed similarly to the THUCTC dataset [2]. The dataset is collected from VnExpress and is extracted for classification tasks. It contains 596,524 samples, each of which consists of five fields: id (index), title, sapo (summary), content (the contents of articles), label (article topic).
@@ 3.4.2 Modified Vietnamese Alpaca
This dataset has been developed using a similar method to that of the Standard Vietnamese Alpaca, with a notable difference in format. While the Standard Vietnamese Alpaca adopts the “instructions/inputs/outputs” format as suggested by [3], our method, influenced by [5], utilizes the “input/output” format. This alteration allows for the generation of more diverse samples and more lengthy output, employing GPT-4, GPT-3.5, and GPT-3.5-instruct. The resulting Modified Vietnamese Alpaca dataset contains 50K samples.
@@ 3.5 Vietnamese Self-chat dataset
This dataset is designed to enhance the model’s ability to engage in multi-turn conversations with humans. We follow two steps:
@@ 3.6 Vietnamese Roleplay Realm dataset
This is a dataset of GPT-generated characters made to increase the ability of open-source language models to role-play. It contains 446 characters generated by GPT-3.5. To construct this dataset, we follow four steps:
@@
  • Topic Generation: We then created conversation topics for each character, drawing from their descriptions. The output field for this step is ”topics”. We generates 20 topics for each character.
  • Dialogue generation: Based on the character descriptions and topics, we generated dialogues. The output for this step is encapsulated in the ”dialogues” field.
@@ 5 Acknowledgments
This work was funded by NAVER Corporation.

### 05 Mar 2024 v2

@@ 2 Main contributions
  • A massive Vietnamese NewsCorpus dataset of around 32M articles, a substantial 53 GB in size, rigorously cleaned, deduplicated, and formatted specifically for the continual pretraining of LLMs.
@@
  • An additional large-scale Vietnamese NewsCategory dataset in a “text-category” format, specifically designed for text classification tasks.
  • Vietnamese Alpaca datasets, tailored for supervised fine-tuning LLMs.
@@
  • We also offer two base models, Vietnamese LLaMA2-7b, which have been further pretrained on an expansive corpus of Vietnamese text, 40 GB and 120 GB, respectively, derived from LLaMA2, marking a significant advancement in the understanding and generation of the Vietnamese language.
@@ 3.2 Vietnamese NewsSapo dataset
The Vietnamese NewsSapo dataset was constructed to train sentence/passage embeddings. Our dataset is structured in a ”title-abstract-contents” format, where each news article is represented by a tuple of (title, abstract, content). The content is the main text body of the article and has been processed to remove images, videos, and other non-textual elements. The dataset contains 31,728,183 triples.
@@ 3.3 Vietnamese NewsCategory dataset
The Vietnamese NewsCategory dataset is constructed similarly to the THUCTC dataset [2]. The dataset is collected from VnExpress and is extracted for classification tasks. It contains 596,524 samples, each of which consists of five fields: id (index), title, sapo (summary), content (the contents of articles), and label (article topic).
@@ 3.4.2 Modified Vietnamese Alpaca
This dataset has been developed using a similar method to that of the Standard Vietnamese Alpaca, with a notable difference in format. While the Standard Vietnamese Alpaca adopts the “instructions/inputs/outputs” format as suggested by [3], our method, influenced by [5], utilizes the “input/output” format. This alteration allows for the generation of more diverse samples and more lengthy output, employing GPT-4, GPT-3.5, and GPT-3.5-instruct. The resulting Modified Vietnamese Alpaca dataset contains 25K samples.
@@ 3.5 Vietnamese Self-chat dataset
This dataset contains around 30K dialogues designed to enhance the model’s ability to engage in multi-turn conversations with humans. We follow two steps:
@@ 3.6 Vietnamese Roleplay Realm dataset
This is a dataset of GPT-generated characters made to increase the ability of open-source language models to roleplay. It contains 446 characters generated by GPT-3.5. The total number of dialogues is about 9K.
To construct this dataset, we follow four steps:
@@
  • Topic Generation: We then created conversation topics for each character, drawing from their descriptions. The output field for this step is “topics”. We generate 20 topics for each character.
  • Dialogue generation: We generated dialogues Based on the character descriptions and topics. The output for this step is encapsulated in the “dialogues” field.
@@ 5 Acknowledgments
This work was funded by NAVER Corporation. We extend our gratitude to PHPC - Phenikaa University and NVIDIA for their generous provision of computing resources for model training.